If a work cell had the following data, what is their OEE or overall equipment effectiveness?

2 – 8 hour shifts

30 minute break per shift

2 hours planned downtime per shift

1 - 30 minute meeting per shift

Cycle time = .025 hours/unit

350 units were produced

325 were good parts
Answer #1

The Overall Equipment Effectiveness is given by following formula:

OEE = Availability (A) x Performance (P) x Quality (Q)

1.

Availability = Run Time/Planned Production time

Run time = Planned Production time – Losses – Breaks

Planned Production time = Shift length – Breaks = (2 shifts x 8 hours per shift) – (2 shifts x 30 minutes breaks/shift) = 16 hours – 1 hour = 15 hours

Planned Production time (in minutes) = 15 hours per day

Run time = 960 mins – (time loss due to (downtime + meeting))

2 hours planned downtime per shift

1 - 30 minute meeting per shift

Run time = 15 – (2*2 + 2*0.5) = 10 hours

Availability = Run Time/Planned Production time = 10/15 = 0.6667

2.

Performance = (Ideal cycle time x total count)/Run time

Cycle time = 0.025 hours/unit

Total count = 350 units were produced

Performance = (0.025 hrs/unit x 350)/10 = 8.75/10 = 0.875

3.

Quality = Good Count/Total Count

Good count of parts = 325 units

Total count = 350 units

Quality = 325/350 = 0.9285

OEE = A x P x Q = 0.6667 x 0.875 x 0.9285 = 0.5416

OEE = 54.6%
Answer #2

To calculate the Overall Equipment Effectiveness (OEE), we'll use the formula:

=××

where:

  • Availability = (Total Operating Time - Planned Downtime) / Total Operating Time

  • Performance = (Total Count \times Ideal Cycle Time) / Total Operating Time

  • Quality = Good Count / Total Count

Given the following data:

  • Total Operating Time = 2 shifts × 8 hours/shift = 16 hours

  • Planned Downtime = 2 hours/shift × 2 shifts = 4 hours

  • Total Count = 350 units

  • Good Count = 325 units

  • Cycle Time = 0.025 hours/unit

Now, let's calculate the components of OEE:

  1. Availability: Availability = (Total Operating Time - Planned Downtime) / Total Operating Time Availability = (16 hours - 4 hours) / 16 hours Availability = 12 hours / 16 hours Availability ≈ 0.75

  2. Performance: Performance = (Total Count × Ideal Cycle Time) / Total Operating Time Performance = (350 units × 0.025 hours/unit) / 16 hours Performance = 8.75 hours / 16 hours Performance ≈ 0.547

  3. Quality: Quality = Good Count / Total Count Quality = 325 units / 350 units Quality ≈ 0.929

Now, let's calculate the OEE:

OEE = Availability × Performance × Quality OEE ≈ 0.75 × 0.547 × 0.929 OEE ≈ 0.37

The Overall Equipment Effectiveness (OEE) for the work cell is approximately 0.37, which represents 37%.

answered by: Hydra Master
Answer #3

To calculate the Overall Equipment Effectiveness (OEE), we need to consider three main factors: Availability, Performance, and Quality. OEE is the product of these three factors and provides a measure of how effectively the equipment is being utilized.

  1. Availability: Availability measures the percentage of time the equipment is available for production. It is calculated as the ratio of operating time to planned production time.

Operating Time = Total shift time - Breaks - Planned downtime - Meeting time Operating Time = 2 shifts * (8 hours/shift) - 2 * (0.5 hours) - 2 * (2 hours) - 2 * (0.5 hours) Operating Time = 16 hours - 1 hour - 4 hours - 1 hour Operating Time = 10 hours

Planned Production Time = Total shift time - Breaks Planned Production Time = 2 shifts * (8 hours/shift) - 2 * (0.5 hours) Planned Production Time = 16 hours - 1 hour Planned Production Time = 15 hours

Availability = (Operating Time / Planned Production Time) * 100 Availability = (10 hours / 15 hours) * 100 Availability = 66.67%

  1. Performance: Performance measures the actual production rate compared to the ideal production rate. It takes into account any speed losses or inefficiencies during production.

Ideal Production Rate = 1 / Cycle time Ideal Production Rate = 1 / 0.025 hours/unit Ideal Production Rate = 40 units/hour

Actual Production Rate = Total units produced / Operating Time Actual Production Rate = 350 units / 10 hours Actual Production Rate = 35 units/hour

Performance = (Actual Production Rate / Ideal Production Rate) * 100 Performance = (35 units/hour / 40 units/hour) * 100 Performance = 87.5%

  1. Quality: Quality measures the percentage of good parts produced compared to the total parts produced.

Quality = (Good Parts / Total Parts Produced) * 100 Quality = (325 units / 350 units) * 100 Quality = 92.86%

Now, we can calculate the Overall Equipment Effectiveness (OEE):

OEE = Availability * Performance * Quality OEE = 0.6667 * 0.875 * 0.9286 OEE ≈ 0.5614 or 56.14%

The Overall Equipment Effectiveness (OEE) for the work cell is approximately 56.14%. This indicates that the equipment is operating at about 56.14% of its full potential efficiency, considering factors such as availability, performance, and quality.


answered by: mervetokaz